# Eaman Jahani

### v2

@@ Abstract
In an online experiment with N = 1,893 participants, we collected and analyzed over 18,000 prompts and over 300,000 images to explore how the importance of prompting will change as the capabilities of generative AI models continue to improve. Each participant in our experiment was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor DALL-E 3, or a version of DALL-E 3 with automatic prompt revision. Participants were then asked to write prompts to reproduce a target image as closely as possible in 10 consecutive tries. We find that task performance was higher for participants using DALL-E 3 than for those using DALL-E 2. This performance gap corresponds to a noticeable difference in the similarity of participants’ images to their target images, and was caused in equal measure by: (1) the increased technical capabilities of DALL-E 3, and (2) endogenous changes in participants’ prompting in response to these increased capabilities. More specifically, despite being blind to the model they were assigned, participants assigned to DALL-E 3 wrote longer prompts that were more semantically similar to each other and contained a greater number of descriptive words. Furthermore, while participants assigned to DALL-E 3 with prompt revision still outperformed those assigned to DALL-E 2, automatic prompt revision reduced the benefits of using DALL-E 3 by 58%. Taken together, our results suggest that as models continue to progress, people will continue to adapt their prompts to take advantage of new models’ capabilities.
The release of ChatGPT in November 2022 has fueled a surge of interest in generative AI and large language models across industry, the public sector, and academia, with early research suggesting that generative AI may be a general purpose technology (Eloundou et al., 2024). Indeed, generative AI models have already been found to outperform humans on text annotation tasks (Gilardi et al., 2023) and improve the productivity of workers across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023). However, the extent to which generative AI can provide value is contingent on the quality of the textual instructions, or “prompts,” that users write to a model. As generative AI has proliferated, the practice of creating and iterating on such prompts, often using a variety of specialized techniques (e.g., chain-of-thought and few-shot prompting) (Schulhoff et al., 2024), has come to be known as prompt engineering. Some firms now even offer six-figure salaries for dedicated prompt engineers (Quilty-Harper, 2023), and an expanding academic literature focuses exclusively on prompting-related topics (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023a; Xie et al., 2023). Despite the growing interest in prompting, some argue that as models become more capable at transforming human language into output, the need to craft prompts in specific, and sometimes esoteric, ways will diminish (Acar, 2023; Mollick, 2023). If this comes to pass, the benefits of generative AI will be accessible to a much larger population.
@@
In this paper, we bring experimental evidence to bear on this debate by presenting results from a pre-registered online lab experiment with $N=1,893{}$ participants recruited on Prolific. [^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration detailed in the SI. In the experiment, participants were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”). The LLM-based prompt revisions in the third condition were not made visible to participants during the experiment. [^2] Automatic prompt revision is the default behavior of the DALL-E 3 API endpoint. We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts. While this drastically reduced the number and extent of LLM-based prompt revisions, it did not completely eliminate them. Therefore, the “DALL-E 3” treatment arm still includes some automatic prompt revisions. For brevity and simplicity, we refer to this treatment arm as “DALL-E 3” throughout the paper. Full details on the system prompt’s efficacy are provided in the SI. The task required each participant to make at least 10 attempts at recreating a “target image” as closely as possible by prompting their assigned model. Each participant’s target image was randomly selected from a curated set of 15 images; this set of images was constructed to span realistic use cases of AI-generated images. Throughout the task, participants could view all of their past prompts and generated images alongside the target image. Participants were paid \$4 USD to complete the task, and to incentivize performance, they received \$8 USD if they were in the top 20% of participants, as measured by the similarity of their best attempt to the target image.
@@
Results
The majority of our analyses focus on the DALL-E 2 and DALL-E 3 treatment arms, as they are more directly comparable. We first estimate the effect of access to a more advanced generative model on task performance. [^3] All main text analyses use ”replay” data from Figure 2 for consistency and to avoid model drift issues. All results hold with the “original” data. See the SI for details. Given the stochastic nature of generative AI models, our measure of task performance is the expected similarity of images generated from each prompt to the relevant target image. To construct a quantitative measure of this expected image similarity for each prompt, we first used each participant-generated prompt to produce 10 images. We then computed CLIP embedding vectors (Radford et al., 2021) for both target images and prompt-generated images. For each participant prompt and its corresponding target image, we then calculated the average cosine similarity ($CoSim$) between the embedding vectors of the randomly sampled generated images and the target image vector (see SI for further details on how we quantify expected image similarity). Unsurprisingly, we find that providing access to a more advanced generative model improved task performance. Participants using DALL-E 3 produced images that were, on average, $z=0.19$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$) than those produced by participants using DALL-E 2. Figure 1A presents three examples from our data to illustrate the qualitative magnitude of this average treatment effect (ATE), which corresponds to a considerable increase in similarity to the target image. We also find that the treatment effect increased slightly as participants made successive attempts to replicate the target image (See Figure 1B; $\beta=0.0010$, $p=0.0227$, see SI for full regression). In other words, the performance gap widened between DALL-E 2 and DALL-E 3 as participants gained more experience with the respective models, potentially due to changes in prompting.
@@ Results
To investigate how participants’ prompts changed in response to DALL-E 3’s more advanced capabilities, we conducted an exploratory analysis of the contents of participant prompts and how those prompts evolved as participants iterated. We find that participants assigned DALL-E 3 wrote prompts that were 24% longer ($\Delta Words=6.9$, $p<10^{-5}$), and that the difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-8}$). Despite differing in length, prompts written for both models contained similar proportions of nouns and adjectives (48% DALL-E 3 vs. 49% DALL-E 2; $p$ = $0.215$). This suggests that prompts submitted to DALL-E 3 were longer because they conveyed additional descriptive information, not because they were padded with superfluous words. We also find that participants blindly assigned to DALL-E 3 wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0184$, $p=0.02$) and in the aggregate ($\beta=0.0191$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses). These findings indicate that exposure to different models did in fact induce differences in prompting, and that these differences grew over successive replication attempts.
To understand the extent to which the ATE we observe is attributable to these behavioral differences, as opposed to DALL-E 3’s more advanced capabilities, we conducted another exploratory analysis in which we regenerated images with (or “replayed”) all participant prompts on both DALL-E 2 and DALL-E 3, regardless of the model the participant used during the experiment. This allows us to decompose the ATE into two components: the “model effect,” i.e., the average increase in similarity to the target when replaying the prompts written by DALL-E 2-assigned participants on DALL-E 3, and the “prompting effect,” i.e., the difference between the average similarity to the target of all prompts generated by DALL-E 3-assigned participants and the average similarity to the target of all prompts generated by DALL-E 2-assigned participants, both replayed on DALL-E 3. Figure 2 shows that the model effect (red) accounts for 51% of the ATE ($\Delta CoSim=0.00841$, $p<10^{-8}$), whereas the prompting effect (blue) accounts for 48% of the ATE ($\Delta CoSim=0.00788$, $p=0.024$). Thus, nearly half of the ATE is driven by endogenous changes in prompting strategy that emerged after exposure to DALL-E 3. The rightmost point in Figure 2 shows that these endogenous changes in prompting took advantage of the unique capabilities of DALL-E 3; when replaying both sets of prompts on DALL-E 2, the difference in performance is not statistically significant and is close to zero ($\Delta CoSim=0.0020$, $p=0.56$). In other words, DALL-E 3 was more capable than DALL-E 2 at effectively rendering the additional information contained in the longer and more descriptive prompts (Betker et al., 2023), and participants adapted their prompting behavior accordingly.
@@
One might think that the need to adapt one’s prompts to the capabilities of different generative models can be obviated by LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023). We use data from our third treatment arm—DALL-E 3 with revision—to test this hypothesis. We find that although participants assigned to DALL-E 3 with revisions performed better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$), prompt revision reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]). Importantly, the positive impact of access to DALL-E 3 with revision was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2). In other words, prompt revision actually caused participants to perform worse in our context. This finding suggests that as currently implemented, AI-assisted prompt revisions are not a panacea, and can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals.
@@ Discussion
In summary, we find that the progression of generative AI models from DALL-E 2 to DALL-E 3 caused performance on our image replication task to increase. This was not only because of DALL-E 3’s superior capabilities (the model effect), but also because of changes in prompting behavior that took advantage of those capabilities (the prompting effect). More specifically, participants using DALL-E 3 wrote longer prompts that contained more descriptive information, and DALL-E 3 was able to render that additional information more effectively than DALL-E 2. These changes in prompting emerged despite the fact that participants were unaware of which model they were assigned. Although automated LLM-revision features are designed to and have been shown to help people generate higher-quality content (Betker et al., 2023; Li et al., 2024), we also find that LLM-based prompt revision failed to further improve performance; instead, automatic prompt revision reduced the benefits of using DALL-E 3. Overall, we find that shifting to a more capable model, whether or not accompanied by automatic LLM revisions, did not eliminate the need for prompting.
The primary limitation of our work is that we only studied the transition from DALL-E 2 to DALL-E 3. We leave the study of future transitions and different types of generative AI models to future work. However, if the trends we report generalize and continue into the future, one could imagine a lock-step dynamic where as models continually improve, people continually respond by adapting their prompts to take advantage of the newest model’s capabilities. Such a pattern suggests that as generative AI models advance, newer models are unlikely to render prompting obsolete. Rather, prompting will be the mechanism by which people unlock new models’ capabilities. Thus, we anticipate that prompting will remain relevant in the future, and that it will be an important determinant of who is best positioned to use and directly benefit from generative AI moving forward.
@@ B.5.4 Successive Prompt Token Sort Ratio
Starting with each participant’s second prompt, we also calculated the token sort ratio (TSR) of each prompt $p_{i,n}$ to the immediately preceding prompt $p_{i,n-1}$. TSR is a fuzzy string-matching technique (Singla and Garg, 2012b) that provides a continuous measure of how similar two strings are. We refer the reader to (Bosker, 2021) for a more in-depth description of how TSR is calculated.
@@ B.5.6 Prompt Composition
We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt. SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation posuniversaldep. These tags encompass parts of speech such as adjectives, adverbs, nouns, and verbs. The model tags each word in a prompt according to this framework, after which we count the total number of words corresponding to each part of speech for each prompt.
@@ Appendix F Pre-registration
Below, we provide a high-level description of all of our pre-pregistered analyses, along with a description of the results of those analyses. In our pre-registration, we declared the intent to conduct each of our analyses using six777see section F.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables. possible outcome variables, all of which are different transformations of the same underlying data: CLIP embedding cosine similarity and DreamSim, both of which rescaled in three ways:
@@ F.1 Hypotheses and Results
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores, as opposed to task-iteration Z-scores). When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks). When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
@@
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores). When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find the same pattern of head-to-head statistical significance results. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks). When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.

### v3

@@ Abstract
The recent surge in generative AI has led to new models being introduced almost every month. In light of this rapid progression, we pose and address a central question: to what extent must prompts evolve as the capabilities of generative AI models advance? To answer this question, we conducted an online experiment with N = 1,893 participants where each participant was incentivized to write prompts to reproduce a target image as closely as possible in 10 consecutive tries. Each participant was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor, DALL-E 3, or a version of DALL-E 3 with automatic prompt revision. In total, we collected and analyzed over 18,000 prompts and over 300,000 images. We find that task performance was higher for participants using DALL-E 3 than for those using DALL-E 2. This performance gap corresponds to a noticeable difference in the similarity of participants’ images to their target images, and was caused in equal measure by: (1) the increased technical capabilities of DALL-E 3, and (2) endogenous changes in participants’ prompting in response to these increased capabilities. Furthermore, while participants assigned to DALL-E 3 with prompt revision still outperformed those assigned to DALL-E 2, automatic prompt revision reduced the benefits of using DALL-E 3 by 58%. Our results suggest that for generative AI to realize its full impact on the global economy, people, firms, and institutions will need to update their prompts in response to new models. Not doing so could leave more than half of the potential benefits of these AI systems untapped.
Recent economic forecasts predict that generative AI could add trillions of dollars to the global economy annually, and there is mounting evidence that this new technology is being integrated into work practices in areas as diverse as business, medicine, and government FactSet (2024); Zhang and Kamel Boulos (2023); Bright et al. (2024). This increasing generative AI adoption appears to be warranted, with recent studies finding that the adoption of generative AI improves the productivity of workers across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023). A common feature of these studies is that the model participants interacted with did not change; however, generative models are continuously updated, with new versions being released as often as once a month. Humans interface with generative models by providing textual instructions or “prompts” to the models (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023; Xie et al., 2023). It remains an open question whether organizations will need to adapt these prompts as models improve to realize the full economic and societal benefits of generative AI. For example, it is now common amongst software companies to have a hundred or more development teams, each with dozens, if not hundreds, of prompts integrated with their code. On the one hand, firms might write these prompts once and subsequently benefit from new models with minimal additional effort. Alternatively, as models improve, organizations may need to continually adapt their prompts. In this paper, we investigate whether this prompt adaptation process is necessary, and the extent to which the ability to adapt one’s prompts is a specialized skill.
@@
We do so by conducting a pre-registered online lab experiment in which $N=1,893{}$ participants from Prolific were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”). [^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration detailed in the SI. The LLM-based prompt revisions in the third condition were not made visible to participants during the experiment.222prompt revision is the default behavior of the DALL-E 3 API endpoint. We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts. While this drastically reduced the number and extent of LLM-based prompt revisions, it did not completely eliminate them. Therefore, the “DALL-E 3” treatment arm still includes some automatic prompt revisions. For brevity and simplicity, we refer to this treatment arm as “DALL-E 3” throughout the paper. Full details on the system prompt’s efficacy are provided in the SI. The task required each participant to make at least 10 attempts at recreating a “target image” as closely as possible by prompting their assigned model. Each participant’s target image was randomly selected from a curated set of 15 images; this set of images was constructed to span realistic use cases of AI-generated images. Throughout the task, participants could view all of their past prompts and generated images alongside the target image. Participants were paid \$4 USD to complete the task, and to incentivize performance, they received \$8 USD if they were in the top 20% of participants, as measured by the similarity of their best attempt to the target image.
Results
To understand how organizations might benefit from model improvements, we first examined whether access to a more advanced generative model enhanced task performance. Comparing the DALL-E 2 and DALL-E 3 treatment arms, we find that providing access to a more advanced generative model improved task performance. Participants using DALL-E 3 produced images that were, on average, $z=0.19$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$), and the size of this gap did not shrink as participants gained more experience with the task (See Figure 1B; $\beta=0.0010$, $p=0.0227$, see SI for full regression). Figure 1A presents three examples from our data to illustrate the qualitative magnitude of this average treatment effect (ATE), which corresponds to a considerable increase in qualitative similarity to the target image.
@@ Results
Access to a more advanced generative model not only increased task performance but also induced changes in the way that participants wrote prompts, which we highlight through two exploratory analyses. First, prompts to the more advanced model were longer and contained more information. Those assigned to DALL-E 3 wrote prompts that were 24% longer ($\Delta Words=6.9$, $p<10^{-5}$), and this difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-8}$). Despite this difference in length, prompts written for both models contained similar proportions of nouns and adjectives (48% DALL-E 3 vs. 49% DALL-E 2; $p$ = $0.215$), suggesting that prompts submitted to DALL-E 3 conveyed additional descriptive information. Second, those assigned to DALL-E 3 explored the space of possible prompts differently; they wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0184$, $p=0.02$) and in the aggregate ($\beta=0.0191$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses). In other words, those assigned to DALL-E 3 were more likely to thoroughly explore a smaller set of prompting approaches.
The changes in prompting we observe are an important mechanism through which participant performance improves, suggesting that organizations will need to adapt their prompts as models evolve. More specifically, using the decomposition procedure described in the Materials and Methods, we find that changes in prompting accounted for 48% of the performance ATE ($\Delta CoSim=0.00788$, $p=0.024$; blue in Figure 2), whereas the shift to a more capable model accounted for 51% of the performance ATE ($\Delta CoSim=0.00841$, $p<10^{-8}$; red in Figure 2).
We also supplied the prompts written for DALL-E 3 to DALL-E 2 (see Materials and Methods for details). Doing so did not improve performance ($\Delta CoSim=0.0020$, $p=0.56$), indicating that DALL-E 3 prompts specifically took advantage of its superior ability to render information (Betker et al., 2023).
Our results show that adapting prompts is crucial for maximizing the benefits of more capable generative models, raising the question of how organizations should approach this adaptation process. One possibility is that LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023) obviates the need for humans to tailor their prompts to different generative models. Comparison of our third treatment arm–DALL-E 3 with revision–to the DALL-E 3 arm highlights the potential pitfalls of this approach. We find that prompt revision actually caused participants to perform worse in our context as it reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]).
@@
Furthermore, the positive impact of access to DALL-E 3 with revision performed only slightly better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$) and was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2). These findings suggest that as currently implemented, AI-assisted prompt revisions are not a panacea. They can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals.
[…]
@@ Discussion
The primary limitation of our work is that we only studied the transition from DALL-E 2 to DALL-E 3. We leave the study of future transitions and different types of generative AI models to future work. Should our results generalize to other models, then we would predict that firms that do not update their prompts as models progress are leaving roughly half the benefit of AI unrealized. Software systems, models, and prompts are becoming increasingly interwoven which is a trend we expect to accelerate. Thus, the lack of prompt maintenance could, in turn, dampen the economic impact of AI on the global economy. Fortunately, updating prompts can be learned by humans across the skill distribution even if models are not currently able to update prompts in an automated fashion. Thus, we anticipate that as firms adopt generative AI, they will need to regularly refine prompts to fully realize the potential of evolving models. In particular, one could imagine a lock-step dynamic where, as models continually improve, people will need to—and are easily able to—respond by adapting their prompts to take advantage of the newest model’s capabilities. Such a pattern suggests that as generative AI models advance, technical prompting infrastructure will not be a one-time investment. Rather, prompting will be the mechanism by which people and firms unlock new models’ capabilities.
Table 1: Model and Prompting Effect by Performance
Effect Estimate p-value
Model $0.011000$ $p<0.00001$
$(0.001860)$
Model $\times$ $-0.000060$ $0.0210$
Performance Decile $(0.000030)$
Prompting $0.010900$ $0.0006$
$(0.003160)$
Prompting $\times$ $-0.000060$ $0.2444$
Performance Decile $(0.000050)$
Material and Methods
Outcome measure Given the stochastic nature of generative AI models, our outcome measure of task performance is the expected similarity of images generated from each prompt to the corresponding target image. To compute this, for each prompt we generated 10 images and computed CLIP embeddings (Radford et al., 2021) for both these and the target images. Each prompt’s expected similarity score was then computed as the mean cosine similarity ($CoSim$) between its generated image embeddings and the target image embedding (see SI for details).
Effect Decomposition We conducted an exploratory analysis where we regenerated images with all participant prompts on both models. This allows us to decompose the ATE into two components: the “model effect” (the average improvement when running DALL-E 2 prompts through DALL-E 3) and the “prompting effect” (the difference in target similarity between DALL-E 3 and DALL-E 2 participants’ prompts, both evaluated on DALL-E 3). Note that all main text analyses use “replay” data from Figure 2 for consistency and to avoid model drift issues. All results hold with the “original” data. See the SI for details.
Experimental Materials Preregistration, data, and analysis code will be deposited at https://osf.io/ejbtp. Full details regarding our experiment design, analysis techniques, and additional results can be found in the SI.
@@ B.5.4 Successive Prompt Token Sort Ratio
Starting with each participant’s second prompt, we also calculated the token sort ratio (TSR) of each prompt $p_{i,n}$ to the immediately preceding prompt $p_{i,n-1}$. TSR is a fuzzy string-matching technique (Singla and Garg, 2012) that provides a continuous measure of how similar two strings are. We refer the reader to (Bosker, 2021) for a more in-depth description of how TSR is calculated.
@@ B.5.6 Prompt Composition
We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt. SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation Universal POS tags (2024). These tags encompass parts of speech such as adjectives, adverbs, nouns, and verbs. The model tags each word in a prompt according to this framework, after which we count the total number of words corresponding to each part of speech for each prompt.
@@ Appendix F Pre-registration
Below, we provide a high-level description of all of our pre-pregistered analyses, along with a description of the results of those analyses. In our pre-registration, we declared the intent to conduct each of our analyses using six666see section F.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables. possible outcome variables, all of which are different transformations of the same underlying data: CLIP embedding cosine similarity and DreamSim, both of which rescaled in three ways:
@@ F.1 Hypotheses and Results
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores, as opposed to task-iteration Z-scores). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
@@
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find the same pattern of head-to-head statistical significance results. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.